Technical Observing for 3-4 September 2008 MST

Observers: D. Miller, A. Rakich and J. Hill in Tucson
Telescope Operator: A. Ceranski

Summary

Lost the first 2 hours to weather, then lost 2 hours to a fuse problem in the E-stop system. We got 2 hours to adjust pointing and confuse ourselves about rotator angle. Closed dome due to rising humidity.

Night time Activities

Times in UT below

02:30 Residue of thunderstorms in the area so we can't open. However, we make good use of the time by having the LUCIFER team preset the telescope (successfully) with their software.

03:00 Stephen and John repaired and tested the problem with LBC and the new IIF build. The present LBC build can properly connect to TCS.

03:30 Have changed LEFTZEROPOINT to 0.0 and LEFTIAA to -65.9 (a 60 degree change from two nights ago) as they clocked the AGw by 60 degrees on the rotator on Tuesday. Soon we'll see if I guessed the correct sign.

04:00 Weather has cleared so we can open the dome.

04:10 Just after our first preset, the telescope stops dead. HBS is off, both Az and El MCRs have dropped out. It turns out to be a problem with the Emergency Stop system.

06:00 John L. and Aaron have tracked down and replaced the blown fuse in the E-stop system so we are on the sky again. Lost time: 2 hours

06:10 pointo M5_0881 to make the nightly adjustment to the pointing model offset. With the default IE=-56 CA=-37 we need offset 0 80. (Why are these 90 arcsec from where we were the previous nights?, or are they 90 arcsec from the default pointing model?)

06:30 The pointing model offsets are IE=-110 CA=-110 (with slightly imperfect collimation).

06:36 pointto BS_9188

Change LEFTIAA to -185.9, and updated IE=-110 CA=-110

06:46 Restart PCS and GCSL

06:53 Restart PCS again after fixing John's editting mistake in the pointing model. (The same mistake explains a bit of the pointing issues on Monday night.)

06:55 Now we see the star, but Pointto failed to start guiding since Doug was still taking manual images with GCS.

06:57 Pointto failed since rotator wasn't ready.

06:59 Pointto failed because it couldn't find a star. But we just saw the star 2 minutes ago? We found it at +15 -30.

07:14 Restarted PCS with the historical value of IA=-286.9358.

07:15 Pointto BS_9188 again

IE=-115 CA=-65

07:35 Pointto M5_0042 with IRTC

Refined center IE=-107 CA=-56

07:41 back to BS_9188 to try guiding, but the offset fails.

07:49 Change LEFTIAA from -185.9 to -5.9, restart PCS.

07:56 Pointto BS_9188 to try guiding again, but also the offset fails.

08:00 GCS hangs after Doug presses the Stop guiding button. Had to kill and restart GCS.

08:02 Close dome for high humidity.

Mystery #1: Why does the pointing seem to change by 10 - 15 arcsec everytime we preset to a new star? (Based on a handful of stars, the pointing seems worse than it was in May/June.)

Mystery #2: What is the correct value of LEFTIAA? We are unclear about which is the correct rotation to give PA-0. (Everytime we get close to measuring it, something else breaks.)

08:30 Still closed, but sending active optics corrections to PSF in the default setting. The appropriate Z4, Z7, Z8 zernikes go to the secondary mirror, and the others go to the primary mirror. However, the secondary did not move. Was this because the Primary failed to reach tolerance at horizon?

Jesper joins us from Potsdam. He reports that the change in UMAC protocol has removed the big pile of communications errors that we had.

08:50 SX mirror panics while slewing up from horizon. -- JohnHill - 04 Sep 2008
Topic revision: r1 - 04 Sep 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback